OpenStudy (anonymous):

What is the solution of the absolute value of x minus 4 is less than or equal to negative 5 ? A. x ≤ -1 and x ≥ 9 B. x ≤ -1 or x ≥ 9 C. no solution D. all real numbers

OpenStudy (zpupster):

\[\left| x-4 \right| \le -5\]

OpenStudy (anonymous):

x is less than or eaqual to -1

OpenStudy (anonymous):

i get that part its the other part do i switch the sign and turn -5 positive?

OpenStudy (zpupster):

yes solve these 2 cases and that is you answer

OpenStudy (anonymous):

if i t were a positive 5 and i switched it then would i switch the sign?

OpenStudy (zpupster):

remember x-4 could be positive or negative to accomplish this we do it in cases: case1 x-4 >or= 5 case 2 x-4≤ -5

OpenStudy (anonymous):

ohhh ok thanks